OpenStudy (anonymous):

Solve: |2t + 5| = 3 A. {1, 4} B. {-1, -4} C. {-1, 4} D. {1, -4}

OpenStudy (hoblos):

|2t + 5| = 3 2t+5=3 OR 2t+5=-3 t=-1 t=-4 {-1, -4}